Cassandra wants to determine the weight of her two dogs, Mimi and Mitzi. However, neither dog will stand on the scale by herself. Cassandra, Mimi, and Mitzi altogether weigh 175 pounds. Cassandra and Mimi together weigh 143 pounds. Cassandra and Mitzi together weigh 139 pounds. Set up and solve a system of equations to find out how much Cassandra, Mimi, and Mitzi each weigh.
Answer by TimothyLamb(4379)

x = cassandra weight
y = mimi weight
z = mitzi weight
---
x + y + z = 175
x + y = 143
x + z = 139
---
put the system of linear equations into standard form
---
x + y + z = 175
x + y = 143
x + z = 139
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = cassandra weight = 107 pounds
y = mimi weight = 36 pounds
z = mitzi weight = 32 pounds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---